Show that the functions defined below is solution of the Differential Equations or not

a)𝑦=π‘₯+3π‘’βˆ’π‘₯, 𝑦′+𝑦=π‘₯+1,𝑀hπ‘’π‘Ÿπ‘’ 𝑦′=𝑑𝑦 𝑑π‘₯

b)𝑦=2π‘₯2 +𝑒π‘₯ +6π‘₯+7, 𝑦′′ βˆ’3𝑦′ +2𝑦=4π‘₯2 c)𝑦= 1 , (1+π‘₯2)𝑦′′+4π‘₯𝑦′+2𝑦=0

d) 𝑒 = 𝑒3π‘₯ cos 2𝑦 , 𝑒π‘₯π‘₯ βˆ’ 2 𝑒𝑦𝑦 = 17 𝑒 , 𝑒π‘₯π‘₯ = 𝑑2𝑒 𝑑π‘₯2


1
Expert's answer
2021-05-25T15:51:45-0400

(a) "y=x+3e^{-x}"

"\Rightarrow y'=1-3e^{-x}"


Taking LHS-

"y'+y=1-3^{-x}+1+3^{-x}=2\neq" RHS


Given y is not the solution.


(b) "y=2x^2+e^x+6x+7"


"\Rightarrow y'=4x+e^x+6\\[9pt]\Rightarrow y''=4+e^x"


Taking LHS-

"y''-3y'+2y\\=4+e^x-12x-3e^x-18+2x^2+e^x+6x+7\\=-e^x+2x^2-6x+11\neq RHS"


Given y is not the solution.


(c)𝑦= 1 , "(1+\ud835\udc65^2)\ud835\udc66''+4\ud835\udc65\ud835\udc66'+2\ud835\udc66=0"

As, y=1

"\Rightarrow y'=0,y''=0"


Taking LHS-

"(1+𝑥^2)𝑦′′+4𝑥𝑦′+2𝑦\\=(1+x^2)0+4x(0)+2(1)\\=2\neq RHS"


Hence y=1 is not the solution.


(d)

"𝑢 = 𝑒^{3𝑥} cos 2𝑦\\[9pt]\Rightarrow u_x=3e^{3x}cos2y\\[9pt]\Rightarrow u_{xx}=9e^{3x}cos2y"


"\Rightarrow u_y=-2e^{3x}sin2y\\[9pt]\Rightarrow u_{yy}=-4e^{3x}cos2y"


Taking LHS-

"𝑢_{𝑥𝑥} − 2 𝑢_{𝑦𝑦}\\=9e^{3x}cos2y+8e^{3x}cos2y\\=17e^{3x}cos2y=17u=RHS"


Given y is the solution.